It was determined on the following project that the duration for the tasks contained some uncertainty. After more research, the optimistic and pessimistic values were determined. 1. a. Calculate the expected time and variance (using the 95 percent level of confidence) for each activity. b. Determine the critical path and time for completion c. Calculate the probably that the project will finish in 30 days. Calculate the probably of completing the project in 27 days. Other than the critical path, which is the next riskiest path? What is the probability of completing that path in 27 days? Your boss wants to know what time it will take to complete if you are 99% sure (that is, 99% probable). What do you tell your boss? d. e. Activity Predecessor Opt. (days) Normal (days) Pess. (days) 13 10 4 4 1 4 5 C,D,E 8 IK H,l 4 10
Answer #1

a) For each activity, calculate the expected time (a 4m+ b)/6 For each activity, Variance ơ2-(ba),36. Standard deviation of e

D) Other than critical path, the next riskiest path is A-B-C-H-K

Its expected time = 27.

Variance of this path = 5.194

Std deviation = SQRT(5.194) = 2.279

For target time of 27 days, z (27-27.5)/2.279 = -0.2194

Probability = NORMSDIST(-0.2194) = 0.4132

de) For 99% probability, z = NORMSINV(0.99) = 2.3263

Expected time = 27.833 + 2.3263*2.3746 = 33.4 days